For each problem, find the equation of the line tangent to the function at the given point. Your answer should be a slope-form.


1) y = 2x^2 + 16x + 32 at (-3, 2)

2) y = (x-2)^(1/3) at (3,1)

3) y = -x^3 + 3x^2 -4 at (3, -4)

4) y = -3/ x^2 - 4 at (-1, 1)



1
Expert's answer
2022-01-31T14:58:07-0500

1)


"y'=4x+16"

"slope=m=y'(-3)=4(-3)+16=4"

"y-2=4(x-(-3))"

The equation of the line tangent to the function at "(-3, 2)" is


"y=4x+14"

2)


"y'=\dfrac{1}{3(x-2)^{2/3}}"

"slope=m=y'(3)=\dfrac{1}{3(3-2)^{2/3}}=\dfrac{1}{3}"

"y-1=\dfrac{1}{3}(x-3)"

The equation of the line tangent to the function at "(3, 1)" is


"y=\dfrac{1}{3}x"

3)


"y'=-3x^2+6x"

"slope=m=y'(3)=-3(3)^2+6(3)=-9"

"y-(-4)=-9(x-3)"

The equation of the line tangent to the function at "(3, -4)" is


"y=-9x+23"

4)


"y'=\dfrac{6x}{(x^2-4)^{2}}"

"slope=m=y'(-1)=\dfrac{6(-1)}{((-1)^2-4)^{2}}=-\dfrac{2}{3}"

"y-1=-\dfrac{2}{3}(x-(-1))"

The equation of the line tangent to the function at "(-1, 1)" is


"y=-\dfrac{2}{3}x+\dfrac{1}{3}"